OpenStudy (anonymous):

Solve the system? 2x+y+z=9 -x-y+z=1 3x-y+z=9

OpenStudy (anonymous):

(2x+y+z=9 )+(-x-y+z=1),Add we have: x + 2z=10 (2x+y+z=9)+(3x-y+z=9) , Add have: 5x +2z=18 time(-1)(x + 2z=10) = -x-2z=-10 keep: 5x +2z=18 (-x-2z=-10)+(5x +2z=18), Add Have: 4x =8 x = 8/4 x = 2 2+2z=10 z= (10-2)/2 z =4 solve y : 2(2)+y+4=9 y = 9-8 y=1